OpenStudy (anonymous):

any one help please fan and medal Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6).

OpenStudy (anonymous):

if it's parallel it must have the same slope, -3x

OpenStudy (anonymous):

ok sorry do you want the answer choices

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

@maddiegirl can u help

OpenStudy (espex):

Are you familiar with the equation of a line, y=mx+b?

OpenStudy (anonymous):

yeah is that slope intercept

OpenStudy (espex):

Right, then you know that 'm' is the 'slope' when the equation is in that form.

OpenStudy (anonymous):

yeah

OpenStudy (espex):

Two lines that are parallel MUST have the same slope. So in this case, '-3' is your slope that both lines must share.

OpenStudy (anonymous):

ok

OpenStudy (espex):

So to find the new line, all you must do is solve the eqeuation \(y-y_1=m(x-x_1)\)

OpenStudy (espex):

Where \(y_1\) and \(x_1\) are the points on the new line, and 'm' is the slope of your existing line.

OpenStudy (anonymous):

ok so for the y put the2nd y first

OpenStudy (anonymous):

can u go deeper into that please

OpenStudy (espex):

Your 'y' will be a variable as it will represent ANY 'y' on the parallel line, and your '\(y_1\)' will be a point you know. The same applies for 'x' and '\(x_1\)'. Think you can write the starting equation?

OpenStudy (anonymous):

let see the y is the 6 and the x is -1

OpenStudy (espex):

So y-6=.......

OpenStudy (espex):

Remember you're trying to get this to start: \(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

oh y-6=x-(-1)

OpenStudy (espex):

You forgot your 'm'

OpenStudy (anonymous):

oh yeah srry y-6=m[x-(-1)]

OpenStudy (espex):

Exactly, and what is your 'm' in this case?

OpenStudy (anonymous):

slope

OpenStudy (espex):

Yes, but what's its value?

OpenStudy (anonymous):

is it 12

OpenStudy (espex):

Look at your original equation, a parallel line shares the same slope.

OpenStudy (anonymous):

-3

OpenStudy (espex):

Excellent. Now put that all together into the equation and solve for 'y'.

OpenStudy (anonymous):

so y-6=-3[x-(-1)]

OpenStudy (espex):

Yes, exactly.

OpenStudy (anonymous):

how do I solve it

OpenStudy (anonymous):

oh nvr mind thnk you

OpenStudy (espex):

You're welcome. :)

OpenStudy (anonymous):

r u able to help still @eSpeX

OpenStudy (espex):

You solved this one, not much more to do.

OpenStudy (anonymous):

i mean another one

OpenStudy (espex):

Sure.

OpenStudy (anonymous):

thnx ill post it in another one

OpenStudy (espex):

Okay, I'll keep an eye out for a mention. :)